# Transformer Alignment in Large Language Models

### v1

@@ A.2 Dynamical Motivation
The equality of top left and right singular vectors (RA2) suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
@@
where the product is repeated $L$ times. Expanding the brackets shows that $x(t)$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. [2016] which views ResNets as
@@
However, Veit et al. [2016] do not make any assumptions about the alignment of the various $A_{i}$ matrices. The Residual Alignment phenomenon suggests ResNets as implementing the simpler system
@@
A.3 Experimental and Implementation Details
@@
Due to the high memory requirement for computing Residual Jacobians, for experiments involving Residual Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized. Additionally, due to compute restrictions, alignment of singular vectors of Residual Jacobians was computed on a smaller subset of the 1,200 prompts.
A.4 Additional Experimental Results

### v2

@@ A.4 Dynamical Motivation
The equality of top left and right singular vectors suggests that the linearizations form a simple dynamical system that acts on representations. Consider a homogenous linear ODE:
@@
where the product is repeated $L$ times. Expanding the brackets shows that $x(t)$ can be thought of as a collection of many paths of various lengths, due to the binomial identity. This agrees with Veit et al. (2016) which views ResNets as
@@
However, Veit et al. (2016) do not make any assumptions about the alignment of the various $A_{i}$ matrices. The coupling phenomenon suggests ResNets as implementing the simpler system
@@
A.5 Experimental and Implementation Details
@@
Due to the high memory requirement for computing Jacobians, for experiments involving the Jacobians, NVIDIA Quadro RTX 6000 was used additionally for $13B>n_{\text{param}}\geq 7B$ and corresponding models were quantized. Additionally, due to compute restrictions, alignment of singular vectors of Jacobians was computed on a smaller subset of the 1,200 prompts.
A.6 Additional Experimental Results
Figure 15: Hidden Trajectories in LLMs. Principal components of the trajectories of the hidden representations through various LLMs (columns, decreasing in model size, see Table 1) in the prompt: What is the capital of France? The capital is. Top row: all layers. Middle Row: layers in shallower transformer blocks (layers specified above plot). Bottom Row: layers in deeper transformer blocks (layers specified above plot). Trajectories of each input token (last token ‘is’ is plotted in black) are plotted in latent space, visualized with a 2-dimension principal component projection. Representations proceed in distinct outward directions, especially in the second half of transformer blocks (lower row) during which the norm of representations increases, with possible abrupt change in the last layer (outer points in upper row). A clear direction of movement is visible in each token trajectory.
@@ A.6 Additional Experimental Results
Figure 22: Coefficient of variation of layer-wise expodistance. Variation of layer-wise expodistance (Section 3.3) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 4.2) on a suite of untrained LLMs (Appendix A.1).
Figure 23: Various Measurements of Representations. All measurements were made on 100 prompts taken from the WikiText 2 datsets. (Row 1) Norm of hidden representations as a function of layer depth. (Row 2) Line Shape Score (LSS) of the hidden trajectories as a function of layer depth. (Row 3) Mean equidistance of contiguous hidden trajectories as a function of depth. (Row 4) Entropy of logit vectors as a function of depth. Noted in most plots is a line where the behaviour of the measurement drastically changes.
Figure 24: Linearity Emerges with Training. Two plots displaying the evolution of the linearity of the token trajectories through training. (Left) The LSS as a function of training checkpoint for the variants of the Pythia Scaling Suite Biderman et al. (2023). Here, the LSS is measured over each entire prompt. (Right) The mean LSS as a function of layer depth measured at various checkpoints throughout the Pythia 12B model. Here, the LSS is computed on a window of layers of width $11$, centred at the value given by the x-axis.
Figure 32: LSS sorted by LLM parameters.